OpenStudy (anonymous):

A recipe calls for 3 teaspoons of cinnamon for every 4 cups of flour. If you use 6 cups of flour, how much cinnamon would you need?

OpenStudy (potatoes.ramu):

You need to cross multiply

OpenStudy (potatoes.ramu):

3 tsps of cinnamon - 4 cups x tsps of cinnamon - 6 cups x=6*3/4=18/4=4.5